1. Seven more than a number is 11
2. Twelve less than twice a number is 20.
3. Eleven more than the product of x and 3.
4. The quotient of a number and j is six.

Answer by mananth(16946)

1. Seven more than a number is 11
x+7 =11
x=11-7
x=4
2. Twelve less than twice a number is 20.
2x-12 =20
add 12
2x-12+12 = 20+12
2x=32
divide by 2
2x/2 = 30/2
x= 15
3. Eleven more than the product of x and 3.
3x+11
4. The quotient of a number and j is six.
x/j =3

m.ananth@hotmail.ca